OpenStudy (anonymous):

Find the quotient, 4 ^r ÷ 4^ s is it 4^r÷s or is it 4^r-s?

OpenStudy (anonymous):

Second One..

OpenStudy (anonymous):

can you explain why?

OpenStudy (anonymous):

\[\huge \frac{x^a}{x^b} = x^{a-b}\]

OpenStudy (anonymous):

Because you can write it as: \[\large \frac{1}{x^b} = x^{-b}\] So: \[\large x^a \times x^{-b} = x^{a-b}\]